### v2

# Gait-Conditioned Reinforcement Learning with Multi-Phase Curriculum for Humanoid Locomotion

Abstract
We present a unified gait-conditioned reinforcement learning framework that enables humanoid robots to perform standing, walking, running, and smooth transitions within a single recurrent policy. A compact reward routing mechanism dynamically activates gait-specific objectives based on a one-hot gait ID, mitigating reward interference and supporting stable multi-gait learning. Human-inspired reward terms promote biomechanically natural motions, such as straight-knee stance and coordinated arm-leg swing, without requiring motion capture data. A structured curriculum progressively introduces gait complexity and expands command space over multiple phases. In simulation, the policy successfully achieves robust standing, walking, running, and gait transitions. On the real Unitree G1 humanoid, we validate standing, walking, and walk-to-stand transitions, demonstrating stable and coordinated locomotion. This work provides a scalable, reference-free solution toward versatile and naturalistic humanoid control across diverse modes and environments.
I Introduction
Developing natural and efficient locomotion strategies for humanoid robots remains a core challenge in robotics. Unlike manipulators operating in structured environments, humanoids must adapt to dynamic settings where balance, adaptability, and smooth transitions are essential for real-world deployment. Human locomotion exhibits key biomechanical traits—straight-knee support, coordinated anti-phase arm swings, and smooth heel-to-toe contact—which collectively enhance balance, reduce energy expenditure, and regulate angular momentum [1, 2, 3, 4].
Figure 1: Human-like multi-gait locomotion on the Unitree G1 humanoid, including standing, walking, and running. The learned reference-free policy exhibits straight-knee support, coordinated arm-leg motion, and natural transitions without MoCap references.
These traits are not merely stylistic but serve fundamental functions. Rhythmic arm swing, for instance, helps cancel leg-induced angular momentum, improving trunk stability and reducing metabolic cost. While recent reinforcement learning (RL) work shows emergent anti-phase swing as a side effect of energy minimization [5], such methods rarely enforce angular momentum control directly. In contrast, we propose a biomechanics-inspired reward that explicitly penalizes residual angular momentum and promotes phase-symmetric arm-leg coordination—yielding stable and efficient gaits without relying on trajectory references.
Recent RL advances have enabled agile locomotion for legged robots [6]. However, reference-based methods like Adversarial Motion Priors (AMP)  [7] are constrained by dataset coverage and morphology gaps, while reference-free approaches often result in unnatural or inefficient behaviors, especially for running or transitions.
These limitations are amplified when integrating multiple gaits—standing, walking, running—into one controller. Prior approaches rely on multi-policy distillation [8], mixture-of-experts [9], or policy fusion [10], but such architectures introduce complexity, latency, and reward interference.
We propose a simpler alternative: a unified, gait-conditioned RL framework where a single recurrent policy learns standing, walking, running, and smooth transitions. A gait-conditioned reward routing mechanism activates gait-specific rewards based on a compact gait ID in the observation, mitigating interference and supporting stable multi-gait training.
To encourage human-like motion, we incorporate biomechanically grounded reward terms that promote straight-knee support, arm-leg coordination, minimal foot drag, and push-off dynamics. Our structured multi-phase curriculum enables progressive skill acquisition in simulation and robust deployment on hardware. This curriculum is inspired by biological motor development and bio-inspired gait adaptation studies [11]. A visual overview of the learned behaviors is shown in Fig. 1. In summary, our contributions are:
  • A unified, reference-free RL framework for standing, walking, running, and transitions in a single recurrent policy.
  • A gait-conditioned reward routing scheme that mitigates reward interference.
  • Biomechanically grounded reward shaping for efficient, natural locomotion without MoCap.
  • A progressive multi-phase curriculum for skill expansion and stable training.
II Related Work
Learning natural and efficient locomotion has been a longstanding goal in RL for legged robots, especially in bipedal settings where stability and versatility are more demanding [6]. Existing methods can be broadly categorized into reference-based learning, reference-free approaches, modular multi-skill frameworks, and multi-behavior learning via reward or value function decoupling.
II-A Reference-Based Locomotion Learning
Early successes in humanoid locomotion were predominantly achieved through reference-based approaches, where policies are trained to imitate curated motion capture (MoCap) data. A seminal example is AMP [7], which introduced an adversarial learning framework that encourages policies to generate human-like motions without explicit trajectory tracking. AMP has since been extended to quadrupedal locomotion [12], humanoid whole-body control [13], and even cross-morphology skill transfer such as enabling quadrupeds to learn bipedal walking behaviors [14], showcasing its versatility across robot forms.
Despite their effectiveness, reference-based methods inherently depend on large-scale, high-quality MoCap datasets and often struggle to generalize beyond the distribution of collected motions. Morphological mismatches between human demonstrators and robotic platforms—such as differences in limb proportions, joint limits, and mass distribution—can further degrade imitation performance. Additionally, controllers trained on curated motions may fail to adapt to unseen commands or unexpected environmental disturbances.
II-B Reference-Free RL
To overcome these limitations, reference-free approaches have gained increasing attention, focusing on reward-driven policy learning without external demonstrations. In quadrupedal locomotion, notable examples include Learning Quadrupedal Locomotion over Challenging Terrain [15], where Central Pattern Generators (CPG) were incorporated into the control loop to promote stable rhythmic gaits, and Learning Robust Perceptive Locomotion for Quadrupedal Robots in the Wild [16], where Policies Modulating Trajectory Generators were introduced to regularize motion generation under challenging outdoor environments. These works demonstrated that careful reward engineering and structured priors can enable robust locomotion without relying on MoCap data.
In the context of bipedal robots, two main strategies have been explored. Some studies have designed fully reference-free frameworks by shaping periodic composition rewards to encourage stable walking cycles [17]. Others have incorporated predefined motion references, such as gait libraries or hybrid zero dynamics trajectories, into RL frameworks [18, 19], blending imitation and reward-driven optimization. While both approaches have achieved significant progress in sim-to-real transfer for bipedal locomotion, challenges remain in achieving fully natural, adaptable, and robust multi-gait control without relying on handcrafted motion references.
However, extending these ideas to humanoids remains challenging due to increased dynamic instability, richer contact modes, and more complex morphology.
II-C Multi-Skill and Modular Architectures
Another important line of work explores how to enable a single agent to perform multiple skills or behaviors. Two widely used strategies are policy distillation [20] and DAgger (Dataset Aggregation) [21], both of which rely on supervised imitation. For example, Han et al. [22] employ multi-expert distillation to compress several task-specific controllers into a single policy capable of diverse locomotion across complex terrains. In contrast, Zhuang et al. [8] utilize DAgger to aggregate expert trajectories for robust policy training in dynamic parkour environments.
While these approaches achieve strong performance, they typically require substantial expert data generation and careful management of training curricula, which may limit their scalability to unseen behaviors or extended gait sets.
Alternatively, modular architectures have been proposed to manage multiple skills within a unified system. Mixtures of experts (MoE) frameworks [9] dynamically select among sub-policies based on context, achieving flexible behavior composition. Perpetual Humanoid Control [9] demonstrated real-time control across diverse human-like behaviors using an MoE-based controller. Multiplicative Compositional Policies (MCP) [10] introduced a multiplicative skill blending mechanism, enabling flexible recombination of low-level primitives. Although modular frameworks offer strong flexibility, they often introduce significant architectural complexity and require skill-specific supervision or expert pretraining, which can limit scalability.
II-D Reward Routing and Value Function Decoupling
A complementary research direction tackles multi-behavior learning by mitigating reward interference through reward routing or decoupled value functions. MultiCritic Actor Learning [23] proposed maintaining a single actor while assigning separate critic heads for different tasks, significantly improving multi-task RL stability. Related ideas have been explored in locomotion settings as well. For instance, CPG-based or phase-conditioned policies [15, 16] combine rhythmic motion priors with RL to stabilize multi-gait behaviors.
Despite these advances, achieving coherent, natural, and efficient multi-gait transitions within a single lightweight policy—without explicit experts, motion references, or hierarchical planners—remains a substantial challenge.
II-E Our Approach
In contrast, our work proposes a streamlined solution: we directly train a single recurrent policy to perform standing, walking, running, and smooth transitions within a unified, reference-free RL framework. To mitigate reward interference across gaits, we introduce a simple yet effective gait-conditioned reward routing mechanism, based on a compact gait ID encoding appended to the observation space. Furthermore, by incorporating human-inspired reward shaping derived from biomechanical insights, we promote biomechanically plausible, energy-efficient, and naturally coordinated motions without reliance on MoCap datasets, skill experts, or switching architectures.
Compared to prior modular, distillation-based, or reference-driven methods, our framework emphasizes minimal architectural overhead, unified end-to-end training, and strong generalization across diverse locomotion behaviors, offering a scalable pathway toward robust and naturalistic humanoid locomotion control.
III Methodology
III-A Problem Setup
We address the problem of learning unified multi-gait control for humanoid locomotion, encompassing standing, walking, running, and transitions between these modes. The task is formulated as a partially observable Markov decision process (POMDP), defined by a tuple $(s_{t},a_{t},P,r_{t},p_{0},\gamma)$, where $s_{t}$ represents the latent system state at time $t$, $a_{t}$ is the action executed by the agent, $P$ denotes the transition dynamics, $r_{t}$ is the reward function, $p_{0}$ is the initial state distribution, and $\gamma$ is the discount factor. The objective is to learn a policy $\pi_{\theta}$ that maximizes the expected cumulative discounted return:
$J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\right]$ (1)
III-B Policy Architecture
Our policy adopts an asymmetric actor-critic architecture, where the actor and critic are implemented as recurrent neural networks composed of a Long Short-Term Memory (LSTM) [24] module followed by a Multi-Layer Perceptron (MLP). In an asymmetric setup [25], the actor network, designed for deployability, receives only proprioceptive observations, while the critic network is provided with additional privileged information during training to improve value estimation.
The use of an LSTM enables the policy to capture temporal dependencies across gait cycles and handle partial observability, while the MLP layers further process the temporal embeddings to output control actions or value estimates. This recurrent structure is particularly important for legged locomotion, where delayed contact effects, balance recovery, and multi-gait transitions demand memory beyond instantaneous observations.
Specifically, the actor observes proprioceptive inputs comprising the base angular velocity and gravity orientation projected into the robot’s local frame, commanded linear and angular velocities, joint position offsets from a nominal default pose, joint velocities, and the previous action taken by the policy. To enable explicit gait-mode awareness, the observation also includes a sine-cosine encoding of the gait phase and a one-hot gait ID vector indicating the current locomotion mode (e.g., standing, walking, running, or transition). These observations, which capture instantaneous and intended locomotion context, are fed into an LSTM followed by an MLP head, outputting a 23-dimensional joint position offset vector. The resulting joint targets are tracked by a joint-level proportional-derivative (PD) controller with fixed gains.
In contrast, the critic network receives an extended observation that augments the actor’s inputs with privileged information, which includes foot contact states, ground reaction forces, friction and restitution coefficients at contact points, binary contact indicators at the hip and knee joints, and externally applied disturbance forces along with their application points. This richer information enables the critic network—structured similarly with LSTM and MLP modules—to provide more stable and accurate value estimates, thus facilitating robust policy learning despite partial observability constraints on the actor.
Both actor and critic leverage temporal memory via LSTM structures, enabling reasoning over gait cycles, delayed contact effects, and transitions across gait phases, even in environments with noisy or limited sensory data.
An overview of our proposed gait-conditioned RL framework, highlighting the asymmetric actor-critic structure, gait-conditioned reward routing, and multi-phase curriculum learning strategy, is depicted in Fig. 2. The figure summarizes how gait mode information conditions both actor and critic observations, clarifying the data flow from observations through network processing, reward computation, and final action selection.
Figure 2: Overview of our proposed gait-conditioned RL framework. A recurrent actor receives proprioceptive states with gait encoding. Gait-conditioned reward masks route mode-specific rewards, enabling multi-gait learning.
III-C Gait-Conditioned Reward Routing
To support diverse locomotion behaviors within a unified policy, we introduce a gait-conditioned reward routing mechanism. At each timestep, the system classifies the current gait mode—such as standing, walking, running, or intermediate transitions (e.g., walk-to-stand, run-to-walk)—based on the commanded velocity and dynamic state. This classification is encoded as a one-hot gait ID and appended to the policy observation, conditioning the actor-critic network on the intended gait.
Each reward term is tagged to a specific gait mode or shared across modes. A gait mask is then applied during training to route only the relevant reward components to the learning objective. For instance, push-off dynamics and short contact duration are activated during running, while foot symmetry and stillness bonuses are used in standing. Transition gaits can inherit or interpolate between adjacent modes’ rewards, allowing smooth behavior blending.
As illustrated in Fig. 3, this mechanism enables independent optimization of each gait behavior without interference, while still maintaining a shared policy backbone. It stabilizes multi-gait learning, ensures clean credit assignment, and supports seamless gait switching within a single recurrent controller.
III-D Gait-Conditioned Reward Routing
To enable multi-gait locomotion within a unified policy, we adopt a gait-conditioned reward routing mechanism. At each timestep, a gait mode ID is inferred based on the commanded velocity and the robot’s dynamic state. This ID is encoded as a one-hot vector and appended to the policy’s observation, conditioning the actor-critic network on the intended gait mode.
During training, a gait mask is applied to selectively activate mode-specific reward terms according to the current gait. As shown in Fig. 3, the complete reward vector includes shared components (e.g., task-level tracking and regulation rewards) as well as gait-related terms (e.g., contact patterns or push-off dynamics) specific to walking, running, or standing. The gait mask filters the reward vector such that only the relevant subset contributes to the learning signal.
This design ensures that each locomotion behavior—standing, walking, running, and their transitions—is optimized independently while coexisting within a single recurrent policy. It effectively mitigates reward conflicts, stabilizes multi-gait learning, and enables clean credit assignment across gait modes.
Figure 3: Illustration of gait-conditioned reward routing during three example modes: Run, Walk, and Stand. The unified reward vector includes task-level and regulation rewards (shared across all gaits), as well as gait-specific terms (e.g., contact, push-off, or stability). A gait mask is used to activate only the relevant terms at each timestep. Ellipses (…) indicate omitted modes (e.g., transitions) or rewards.
III-E Curriculum Learning Strategy
Inspired by the natural progression of human motor development—from static balance to walking and eventually running—we design a biologically grounded curriculum to enable stable acquisition of locomotion skills.
Training all gaits simultaneously often leads to reward conflict and unstable exploration. To address this, we adopt a multi-phase curriculum that gradually introduces gait modes, broadens the commanded velocity range, and incrementally activates coordination mechanisms.
The curriculum progresses along three dimensions:
  • Gait complexity: Starting from walking, we progressively add standing, running, and transitions.
  • Commanded velocity: The speed range expands from low-speed walking to high-speed running (up to $4.0$ m/s).
  • Coordination mechanisms: Reward routing, transition blending, and deceleration smoothing are gradually enabled.
We define three training phases :
  • Phase 1: Low-speed walking (002222 m/s), focusing on rhythm, foot clearance, and knee posture.
  • Phase 2: Standing and walk-to-stand transitions with emphasis on double support and postural stability.
  • Phase 3: Running at higher speeds with rewards for push-off, flight phase, and short ground contact.
This progressive curriculum improves training stability, reduces reward interference, and facilitates generalization across locomotion contexts.
III-F Human-Inspired Gait Design and Reward Shaping
While RL can produce stable locomotion behaviors, the resulting motions often appear overly crouched or energetically suboptimal. To encourage more natural and efficient gait patterns, we incorporate biomechanical principles observed in human locomotion directly into the reward design.
Specifically, our approach draws on the following insights:
  • Straight-knee support improves force transmission and reduces muscular effort during stance phases [26].
  • Anti-phase arm-leg coordination mitigates angular momentum buildup and stabilizes the upper body [1, 2].
  • Gait transitions are gradual and biomechanically necessary—humans do not switch abruptly between running, walking, or standing, but instead adopt intermediate steps to reduce momentum and maintain balance. This motivates our inclusion of explicit transition gaits (e.g., walk-to-stand, run-to-walk), which align with observed human motor strategies [27, 28].
These principles are encoded through both gait-specific and shared reward components. Table I summarizes representative terms used to characterize and stabilize each locomotion mode. While the full reward set includes additional objectives (e.g., symmetry, torque minimization), we omit them here for brevity.
TABLE I: Representative Human-Inspired Reward Terms by Gait Mode
Gait Reward Term Description Walking Contact Pattern Encourage phase-aligned foot contacts based on cyclic gait timing. Foot Clearance Promote sufficient foot lift during swing to avoid dragging. Straight Knee Encourage extended knee during stance to improve support efficiency. Running Contact Pattern Encourage alternating single-leg contact and flight phases. Push-Off Dynamics Reward strong vertical and forward velocity during push-off. Short Contact Penalize prolonged stance to promote dynamic running. Standing Contact Pattern Encourage consistent double-foot support for static balance. Base Stability Penalize base and joint motion to maintain upright posture. Transition Contact Pattern Promote correct contact phasing during gait switching. Smooth Deceleration Encourage gradual reduction in velocity to settle into stance.
Key gait-specific rewards used to stabilize and differentiate locomotion behaviors. Additional terms such as symmetry and coordination are used but omitted for brevity.
III-F1 Angular Momentum Compensation via Arm-Leg Coordination
Human arm swing plays a critical role in balancing the whole-body angular momentum during walking. Biomechanical studies [1, 2, 3] have shown that arm motion primarily serves to counteract leg-induced angular momentum—especially in the yaw direction—thereby enhancing trunk stability and reducing energy expenditure.
Inspired by this, we design a reward to promote human-like, dynamically balanced arm swing without relying on trajectory references. The total centroidal angular momentum $\mathbf{L}_{\text{total}}$ is computed via the standard decomposition [29]
To encourage coordinated arm-leg motion, we define the angular momentum reward as:
$R_{\text{momentum}}=\;$ $-\left(L_{\text{total},z}^{2}\right)-0.4\,\left(L_{\text{la},z}-L_{\text{ra},z}\right)^{2}$ (2)
where $L_{\text{la}},L_{\text{ra}}\in\mathbb{R}^{3}$ denote the angular momentum vectors of the left and right arms, and the subscripts $x$ and $z$ refer to pitch and yaw axes.
This reward encourages:
  • Whole-body momentum minimization: Reduces residual angular momentum, particularly in the yaw direction;
  • Anti-phase yaw swing: Promotes alternating arm motion to counteract leg-induced rotation.
III-G Training Platform
Training is performed in Isaac Gym [16], a GPU-accelerated physics simulator supporting parallel simulation of 1000 humanoid environments with NVIDIA RTX 4090 GPU. Policy optimization employs Proximal Policy Optimization (PPO) [30], with recurrent neural networks for both actor and critic, ensuring high sample efficiency and robust policy performance under partial observability.
III-H Domain Randomization
To improve policy robustness and facilitate sim-to-real transfer, we apply domain randomization during training. Several physical parameters are perturbed, including ground friction coefficient, robot base mass, and joint control properties. Additionally, external disturbances are introduced by applying randomized lateral pushes to the robot’s base velocity at fixed intervals.
These perturbations expose the policy to a wide range of dynamics and environmental variations, encouraging generalizable locomotion behaviors that remain stable under uncertainty.
IV Experiment and Results
We evaluate our proposed gait-conditioned RL framework both in simulation (Isaac Gym) and real robot . The policy is trained to perform standing, walking, running, and smooth transitions between these modes using a single recurrent controller. Our experiments demonstrate that the learned policy achieves robust gait switching across varying commanded velocities, while exhibiting human-like motion and energy-efficient, coordinated behaviors—without relying on motion capture references or hierarchical planners.
IV-A Simulation Results
IV-A1 Gait Switching and Velocity Tracking
To evaluate the controller’s adaptability to varying locomotion demands, we command a velocity profile that ramps up and down between 00 and $3$ m/s. Fig. 4 (bottom) shows that the actual forward velocity closely follows the commanded profile across a range of gaits. The top subplot visualizes contact phases along with background gait mode labels.
The policy exhibits stable transitions between standing (green), walking (blue), running (orange), and transitional gaits: Walk-to-Stand (purple) and Run-to-Walk (red). These transitions occur smoothly without abrupt contact shifts, indicating robust internal gait modulation. In particular, the policy slows down naturally from running into walking (R2W), and settles into a balanced standing posture (W2S) without external resets or mode switches.
Figure 4: Top: Support contact classification with background gait labels including Stand, Walk, Run, Walk-to-Stand (W2S), and Run-to-Walk (R2W). Bottom: Commanded vs. actual forward base velocity. Gait transitions are smooth, dynamically consistent, and reflect internal coordination.
IV-A2 Angular Momentum Coordination
Fig. 5 shows the Z-axis angular momentum of legs, arms, and the combined body over time. Throughout different gait modes—including transitions—the policy maintains low total angular momentum by coordinating the motion of arms and legs in anti-phase.
In walking and running, the arm and leg momenta are of opposite sign and similar magnitude, resulting in effective cancellation. During the R2W and W2S transitions, the coordination continues smoothly as the amplitude of each segment adapts to the underlying velocity and contact mode. This demonstrates that the learned behavior is not only dynamically stable, but also biomechanically efficient.
Figure 5: Z-axis angular momentum of legs (blue solid), arms (red dashed), and total (black dash-dot). The policy achieves momentum coordination across standing, walking, running, and transitional gaits (W2S, R2W), with effective cancellation between arm and leg segments.
These results confirm that our gait-conditioned policy learns to: (i) switch naturally between gaits; (ii) track commands robustly; and (iii) maintain biomechanical coordination.
IV-B Real-World Transfer
To evaluate the transferability of our policy from simulation to physical hardware, we deploy the learned controller on a real Unitree G1 humanoid robot. The deployed policy is directly transferred without any additional fine-tuning, sim-to-real adaptation, or dynamics randomization beyond what was already applied during training.
Our system successfully demonstrates stable standing, smooth walk-to-stand transitions, and walking on real hardware. These behaviors remain coherent and robust under moderate command perturbations and exhibit natural whole-body coordination. In particular, the arm-leg coordination and double-foot support in standing mode transfer well, indicating the effectiveness of our human-inspired reward shaping and gait-conditioned training framework.
Figure 6 shows snapshots of the deployed policy performing walking on the physical Unitree G1. The robot maintains biomechanically plausible postures such as extended knees and coordinated arm-leg motion, closely matching the simulated behavior.
Figure 6: Real-world deployment of our learned policy on the Unitree G1 humanoid. The robot demonstrates smooth transitions between walking and standing, with natural limb coordination and upright posture—transferred directly from simulation without any fine-tuning.
Overall, our results validate that the proposed framework generalizes effectively across modalities and environments, with promising performance in real-world humanoid locomotion.
V Conclusion and Future Work
We proposed a unified gait-conditioned RL framework for humanoid locomotion, combining gait-specific reward routing, biomechanically inspired reward shaping, and multi-phase curriculum learning. The resulting recurrent policy enables standing, walking, running, and smooth transitions, all within a single controller and without reliance on motion capture data.
Our method achieves robust and naturalistic multi-gait behaviors in physics-based simulation using a Unitree G1 humanoid. Initial real-world deployment demonstrates successful transfer of standing, walking, and walk-to-stand transitions. We are currently working to extend this to dynamic running behaviors, which are already stable in simulation.
Future directions include scaling to more complex scenarios such as traversal of uneven or deformable terrain, vision-guided locomotion in dynamic environments, and the incorporation of task-conditioned objectives for whole-body behaviors like climbing, object interaction, and human-robot collaboration.